If 277 J is required to change the temperature of 7.11 g of sodium chloride from 275.0 K to 335.0 K, what is the specific heat capacity of sodium chloride?

A. -0.650 J/g×K
B. 1.54 J/g×K
C. 1.18×105 J/g×K
D. 0.650 J/g×K
E. 6.50×10-4 J/g×K
1
Expert's answer
2019-03-22T04:28:30-0400

"q=cm\Delta T"

"\therefore c= \frac{q}{m\Delta T}"

"c=\frac{277 J}{7.11 g \times (335.0 K-275.0 K)} = 0.650 \frac{J}{g \times K}"

Answer: D. "0.650 \frac{J}{g \times K}"